# CREA: A Collaborative Multi-Agent Framework for Creative Content Generation with Diffusion Models

### v1

@@ 4.1 Qualitative Results
4.2 Quantitative Results
Creative Image Editing. For image editing, we compare our method against state-of-the-art techniques, including LEDITS++, InstructPix2Pix, SDEdit, and TurboEdit, across a range of evaluation metrics. To assess how well the edited objects align with the text descriptions, we compute the CLIP score by measuring the similarity between the generated image and its corresponding text prompt. For evaluating diversity, we use the LPIPS [78] score, which quantifies perceptual distance between images. Unlike conventional usage where lower LPIPS indicates better reconstruction, a higher LPIPS in our case signifies stronger, more transformative edits—a desirable trait in creative tasks. Additionally, we compute the DINO [76] score which measures how well the edited image retains key structural and semantic characteristics of the original image, and Vendi [19] score, which quantifies diversity by calculating the Shannon entropy of the eigenvalues of the similarity matrix among the generated images. As shown in Table 1, our method consistently outperforms all baselines across all evaluation metrics, demonstrating its superior ability to generate creative and diverse image edits.
Creative Image Generation. For creative image generation, we compare CREA against ConceptLab, Flux, and SDXL. As shown in Table 1, our method achieves superior results in LPIPS and VENDI scores, while in CLIP scores, it was outperformed by SDXL. We observe that SDXL tends to generate highly colorful images, which aligns with CLIP’s depiction of creativity. However, it produces the lowest LPIPS scores, indicating that its generated images are highly similar to each other, and it also achieves significantly lower VENDI scores compared to CREA and ConceptLab. While ConceptLab and our method achieve comparable results overall, ConceptLab has significantly lower performance CLIP scores, suggesting that its generated images do not align well with the text prompt. This limitation stems from ConceptLab’s reliance on well-defined subcategories—a core design principle of their method. As shown in Fig. 5, ConceptLab fails when a concept lacks sufficient subcategories, such as its inability to accurately generate a couch object. While ConceptLab attains a high VENDI score, this failure highlights its trade-off between diversity and semantic alignment, limiting its effectiveness.
LLM as a Judge. We use a Multi-modal LLM [15] as a judge to provide a scalable and consistent evaluation of creative image edits, as traditional quantitative metrics (e.g., LPIPS or CLIP) fail to capture subjective qualities like originality, expressiveness, and aesthetic appeal. Therefore, we used LLM as a judge and compared methods for several key aspects of creativity, including Originality, Expressiveness, Aesthetic Appeal, Technical Execution, Unexpected Associations, Interpretability and Depth, and Overall Creativity, simulating human-like subjective assessments [12]. As shown in Table 2, our results significantly outperform other methods across all evaluated aspects, demonstrating the effectiveness of our approach in generating more diverse, meaningful, and visually compelling creative edits.
User Study. To assess the creativity of the generated images based on human perception, we conducted a user study where participants evaluated images using three key criteria. Usability measures how accurately the generated image represents the specified object (Q1). Novelty assesses the uniqueness and originality of the image (Q2). For edited images, we evaluate editing consistency, which examines how well the characteristics of the original image are preserved while incorporating creative modifications (Q1). We run our user study on Prolific.com with 50 participants. Participants rated each criterion on a 5-point Likert scale, ranging from 1 (Not at all) to 5 (Very well). The study was conducted across a diverse set of generated and edited images, providing a comparative analysis of our method against existing baselines.
4.3 Additional Experiments
User-Guided Editing and Generation
Our method enables users to guide the creative process according to their preferences (see Fig. 6 (a)). For example, users can specify not only that a creative edit should be applied to a cup but also include additional conditions, such as ‘monster’, to influence the generation. Our approach incorporates these preferences, steering the creative process accordingly. Compared to the baseline Flux + ControlNet approach, where we provided the prompt ‘A creative <condition> <object>’, our results demonstrate greater creativity while still complying to the given conditions.
Personalization To highlight the broader applicability of our approach, we present personalization results using CREA-generated creative images using an off-the-shelf personalization adapter [67]. Fig. 6 (b) demonstrates how generated subjects can be adapted to various contexts or how their styles can be transferred to create new images. Note that due to the limitations in personalization models, some fine-grained details were not able to capture by [67], however the main characteristics are still preserved.
Video Generation To demonstrate the versatility and extensibility of our agentic framework, we extend CREA to creative video generation, where an initial prompt (e.g., ‘a train’) is automatically transformed into creative scenes. We conduct experiments using the CogVideoX [74] model as the generative backbone. Instead of initializing the process with a static creative blueprint, our Creative Director agent generates a structured creative video plan containing key fields: Subject, Action, Setting, Style, and optional Additional Details. This plan is passed to the Prompt Architect, who composes a coherent and high-creativity video prompt via contrastive prompt fusion, similar to our image generation pipeline. The Generative Executor then synthesizes the video using CogVideoX. This evaluation serves as a proof of concept for applying our multi-agent creativity principles beyond static imagery. As shown in Figure 8, our method produces significantly more imaginative and visually engaging results compared to baseline prompts such as “a creative <object>.”
Method | LPIPS-Diversity | Vendi | Creativity (LLM) |
Base | 0.302 0.129 | 3.19 1.49 | 79.67 9.66 |
+Principles | 0.312 0.142 | 3.59 2.46 | 81.41 7.76 |
+Contrastive | 0.391 0.150 | 3.50 1.64 | 82.31 6.98 |
+Self-Enhancement | 0.414 0.157 | 3.70 1.97 | 83.47 6.68 |
Ablation Studies We perform comprehensive ablation studies to analyze the design choices of CREA (see Fig. 7). (a) Ablation on Generalization to other T2I models: Our method extends effectively to different generative models, such as SDXL and DALL-E. (b) Ablation on Parameter Sensitivity: We ablate CFG values (3.5-40) for Flux and the conditioning scale for ControlNet (0.1-0.5) to evaluate their impact. (c) Ablation on Iterative Refinement: We demonstrate the benefits of our method’s refinement process over multiple iterations. (d) Ablation on Prompt Variations: We explore alternative prompts beyond ‘a creative <obj>’ and test negative prompting (e.g., ‘a normal <obj>’, see Fig. 11) to steer the model away from conventional generations. (e) Ablation on Model Components We conduct an ablation study on key components of our framework, including creativity principles, contrastive prompting, and refinement (see Table 3). The base version represents a baseline where the multi-modal LLM model is simply prompted to generate a creative description for a given object.
5 Limitations
While our framework introduces a novel agentic approach to creative image editing and generation, we would like to address a few limitations. Since our framework utilizes a T2I model such as Flux, certain biases inherent to the generative backbone can affect the outputs. For instance, we observe cases where the background is unintentionally darkened based on the object’s semantics (e.g., editing a “lighted box” often results in darker surroundings to emphasize contrast), or where the model hallucinates human figures even when not prompted (see Fig. 7 (e)). In addition, while CREA extends naturally to video generation using models such as CogVideoX, this extension introduces new challenges since attributes like abstraction or surrealism may unintentionally amplify inconsistencies in motion. Nonetheless, our framework represents a significant step toward enabling autonomous and collaborative creativity in AI. By modeling a multi-agent dialogue grounded in creativity principles, CREA pushes beyond traditional prompt-based pipelines to achieve more creative outputs.
6 Conclusion
In this work, we introduce a novel agentic framework for creative image editing and generation, pioneering a disentangled approach that enables greater flexibility and artistic control. By leveraging specialized agents that collaborate to refine and enhance outputs, our method overcomes the limitations of traditional prompt-to-image models, reducing the burden on users while fostering creativity. We demonstrate the versatility of our approach across editing and generation, and highlight its potential for creative video generation. Our findings suggest that agentic frameworks can serve as a powerful foundation for more autonomous and creative AI systems, opening new directions for creative collaboration between humans and AI.
Supplementary Material
[…]
Appendix A Investigating Creativity Types
Investigating Creativity Types We investigate emerging patterns in generated images to identify key attributes—shape, color, and texture—that influence creativity. Our analysis reveals that our method achieves significantly higher scores in these aspects, leading to the generation of more visually diverse and imaginative objects (see Fig. 9). Compared to baseline models, our approach demonstrates a stronger ability to manipulate these factors, reinforcing its effectiveness in producing unique and aesthetically rich generations.
Appendix B Creativity Principles in the Context of AI
B.1 Theoretical Foundations of Creativity
The study of creativity in artificial intelligence (AI) traverses an extensive scholarly terrain, drawing on insights from cognitive psychology, aesthetics, neurobiology, computer science, and beyond. Table 4 in this work identifies several key criteria - Originality, Expressiveness, Aesthetic Appeal, Technical Execution, Unexpected Associations, and Interpretability and Depth that underpin our creativity assessment framework. These criteria are not arbitrary; rather, they reflect foundational theories such as Boden’s categorization of creativity [5], Guilford’s notion of divergent thinking [23], Amabile’s model of social and motivational factors [1], and neuroaesthetic concepts rooted in the works of Ramachandran and Hirstein [54]. These theories collectively provide a robust foundation for understanding creativity as both a human and machine-driven phenomenon. But how does AI actually manifest these traits?
B.2 Computational Creativity: Capabilities, Gaps
By aligning computational methods with these established principles, researchers can better navigate the complexities of what it means for an AI system to be ‘creative’.
Human-centric views on creativity have traditionally emphasized cognitive mechanisms, motivational influences, and cultural contexts. Boden’s theory, for example, delineates creativity into combinational, exploratory, and transformational processes, each capturing an increasingly profound conceptual shift in how ideas are generated and modified.
In AI, combinational creativity can be observed when generative models—particularly Generative Adversarial Networks (GANs) blend stylistic elements from multiple art traditions into hybrid images [22]. While this merging of features may appear novel, it typically does not fundamentally alter the conceptual space from which the models operate, illustrating the difficulties systems still face in achieving transformational creativity. Guilford’s pioneering work on divergent thinking underscored the importance of producing multiple, varied solutions [23]. AI-driven text-generation platforms like GPT-3 and GPT-4 [51] do mimic aspects of divergent thinking by offering an abundance of plausible completions (Brown et al., 2020). Nevertheless, the question persists as to whether these outputs merely recombine learned statistical patterns or truly reflect a leap in imaginative thought.
Beyond cognition alone, Amabile’s componential model highlights how domain-relevant skills, creative processes, and intrinsic motivation intersect to yield creative outputs [1]. This perspective draws attention to the human–machine interface, stressing that creativity often emerges from iterative co-creation rather than algorithmic autonomy. Meanwhile, theories from neuroaesthetics, such as those posited by Berlyne and Ramachandran and Hirstein, elucidate how phenomena like complexity, novelty, uncertainty, and perceptual principles (e.g., ‘peak shift’) influence artistic and aesthetic preferences (Berlyne, 1971; Ramachandran and Hirstein, 1999). Although deep learning algorithms like Style Transfer [21] or StyleGAN [21] may inadvertently exploit some of these principles by amplifying salient features, their reliance on learned patterns rather than explicit aesthetic intent underscores the gap between AI-generated novelty and human interpretive nuance. While these examples show impressive generative power, a critical challenge lies in evaluating whether such outputs are truly creative in the human sense.
B.3 Evaluating AI-Generated Creativity
The Geneplore Model provides a useful lens through which to view many large-scale generative models, outlining how a creative process involves both ‘generation’ (forming preliminary mental representations) and ‘exploration’ (refining or recombining these representations) [16]. Transformer-based architectures, with their attention mechanisms, can be seen as ‘geneplore-like’ when they generate multiple potential sequences and then rank or filter them for coherence [64]. Yet the final outputs typically optimize likelihood or alignment with training data, rather than undergoing any meta-cognitive reflection akin to a human creator’s iterative, intentional revision.
Assessing creativity in AI also demands a balanced approach to formal and subjective metrics. Researchers have developed quantitative measures such as the Frechet Inception Distance (FID) to capture statistical proximity or ‘realism’ [24]. While FID and analogous measures can be informative, they fail to capture nuances of emotional resonance or interpretive richness that many theories consider crucial to truly creative outputs [56]. This motivates a shift from purely statistical benchmarks to human-aligned, multidimensional assessment tools—especially in systems involving co-creation.
B.4 Human-AI Co-Creation in the Future
Although AI models trained on massive datasets produce highly novel outcomes, the iterative nature of human creativity, often involving extensive editing, critique, and context-sensitive judgment, remains only partially replicated in existing machine learning workflows [1]. Systems such as ChatGPT, for instance, enable multi-turn refinements through user inputs [62]. However, they do not spontaneously self-critique or shift creative direction without explicit prompts.
While deep learning has undoubtedly propelled innovation, it seldom delivers truly transformational creativity, as it primarily samples and recombines patterns from existing data rather than fundamentally restructuring conceptual spaces. One promising direction involves placing humans in the loop, harnessing reinforcement learning from human preferences or enabling interactive co-creation scenarios [62]. In such contexts, the interplay of user feedback and algorithmic adaptability can lead to more nuanced, emotionally resonant outcomes. However, this collaboration raises interpretive questions about authorship and creative agency: if the machine contributes significantly to the ideation process, is the creativity emergent from the human artist, the algorithm, or both [43]? These questions underscore that AI creativity is not merely an engineering issue; it is also an epistemological and philosophical inquiry into the nature of creative process itself.
By anchoring AI-based creativity in well-established theoretical frameworks, as we do through the criteria defined in Table 4, researchers and practitioners can adopt a more holistic lens that extends beyond surface-level novelty to capture deeper semantic, aesthetic, and emotional dimensions. While current generative models illustrate significant strides in producing art that appears creative to casual observers, the absence of context-sensitive reflection, domain-specific constraints, and culturally aware interpretive layers hints at the distance yet to be traveled. The path forward likely lies in hybridizing data-driven models with sophisticated symbolic reasoning, integrating iterative self-critique mechanisms, and maintaining a tight coupling with human feedback loops. This multifaceted approach not only enriches the generative potential of AI but also compels us to revisit and refine our own definitions of what it means to create. Through such rigorous, theory-informed methodologies, we can push AI toward outputs that not only surprise but also resonate with the complexities and depths of human creativity. In aligning theory with implementation, we not only improve generative tools—but also challenge and evolve our own understanding of creativity.
Criterion | Related Theory |
Originality (Novelty & Uniqueness) | Boden’s Theory of Creativity, Guilford’s Divergent Thinking |
Expressiveness (Emotional Impact) | Amabile’s Model, Ramachandran & Hirstein’s Laws |
Aesthetic Appeal (Composition & Harmony) | Martindale’s Aesthetic Model, Berlyne’s Aesthetic Theory |
Technical Execution (Craftsmanship & Skill) | Amabile’s Model, AI Creativity Frameworks |
Unexpected Associations (Surprise & Ingenuity) | Geneplore Model, Boden’s Combinational Creativity |
Interpretability & Depth (Exploration Potential) | Ramachandran’s Laws, Finke et al.’s Geneplore Model |
Appendix C LLM-Based Agents
The next-gen AI is famously touted as the ‘Age of Agents’ by many distinguished researchers and for good reasons. Large Language Model (LLM)-based agents represent a significant advancement in AI-driven task automation and decision-making, leveraging the reasoning, contextual awareness, and generative capabilities of LLMs to function as autonomous entities [73]. Unlike traditional rule-based automation or heuristic-driven AI models, LLM-based agents dynamically interpret user input, plan actions, execute complex workflows, and adapt to evolving scenarios. These agents serve as fundamental building blocks for modern AI applications, acting as cognitive intermediaries that integrate natural language understanding (NLU), tool execution, reasoning, and decision-making into a unified system.
LLM agents are intelligent systems that perceive, reason, and act autonomously by leveraging the capabilities of foundation models like GPT-4. As illustrated above, an LLM agent is composed of three core components: Perception, Brain, and Action. The Perception module gathers multimodal inputs (text, images, APIs, sensor data) from the environment, which are then processed by the Brain—the LLM acting as a reasoning core. This brain performs memory retrieval, knowledge grounding, and decision-making to plan suitable responses. Finally, the Action module executes these decisions via natural language outputs, tool usage, or physical actuators. This perception-reasoning-action loop allows LLM agents to operate in dynamic environments, continuously adapting to new tasks and inputs through generalized reasoning and tool-assisted execution.
C.1 Single Agent Systems
A variety of existing LLM-based agent frameworks follow a single-agent paradigm, in which a solitary large language model is augmented with plugins or external capabilities but does not explicitly coordinate with other agents. For example, AutoGPT [61] autonomously pursues user-defined goals by using tools such as a web browser and memory buffer. Similarly, ChatGPT+ offers a premium plan featuring a code interpreter or plugin ecosystem, enabling program execution or curated tool usage to complete more involved tasks [49]. LangChain Agents [33] provide a modular framework that allows developers to define custom “agents” for action selection (e.g., ReAct agents for reasoning and acting), but they remain single-agent in practice, focusing on how one LLM instance selects and executes a sequence of operations. Transformers Agent [28] offers a similar single-agent mechanism, embedding a set of curated tools to be invoked via natural language instructions. While these single-agent approaches illustrate how LLMs can be augmented with external resources and more sophisticated reasoning loops, their lack of agent-to-agent collaboration can limit problem-solving potential in scenarios where multiple specialized roles or iterative debate might be beneficial.
C.2 Multi-Agent Frameworks
Building on these single-agent foundations, recent frameworks employ multiple, concurrently active LLM-based agents to tackle tasks that may require diverse expertise or role specialization. BabyAGI [75] implements a simple multi-agent pipeline wherein distinct agents handle task creation, task prioritization, and sub-task completion. Although this approach showcases the potential of agent specialization, it enforces a predefined, static communication pattern among agents. CAMEL [37] extends multi-agent design by letting agents converse in assigned “roles” (e.g., user, system, developer) and use an inception-prompting technique for autonomous coordination. However, CAMEL currently does not support tool usage such as code execution. Moreover, works on multi-agent debate demonstrate how multiple LLMs can contest or discuss a problem to improve the factual correctness, reasoning depth, or creative divergence of solutions [38] These debate-oriented approaches concentrate on structured dialogues between agents but do not natively incorporate human involvement or external tool usage. Meanwhile, MetaGPT [27] is a specialized application of the multi-agent paradigm directed at software development, demonstrating how distributed GPT “roles” can collaboratively design and refine code bases. Due to the rapid evolution of many open-source repositories, the characterizations offered here reflect the state of these systems at the time of writing; readers interested in a more extensive survey are referred to [36].
C.3 Autogen
AutoGen is an open-source framework developed by Microsoft Research’s AI Frontiers Lab [48], designed to simplify the creation and orchestration of AI agent systems to enable seamless multi-agent collaboration by building modular, programmable agents that interact in structured conversational workflows. At its core, AutoGen provides a unified interface via the ConversableAgent base class, from which various specialized agents inherit—such as the AssistantAgent for code generation or task completion, the UserProxyAgent for human-in-the-loop interactions, and the GroupChatManager for coordinating group discussions. Each agent is configured with a human_input_mode, system messages, tool-use permissions, and execution constraints, making it easy to orchestrate task-specific, autonomous, or collaborative behaviors. The collaborative framework allows agents to share context, ask clarifying questions, or refine outputs in multiple rounds, supporting complex applications like code debugging, multimodal reasoning, and creative generation. This design facilitates flexible experimentation and customization in agent-based research and real-world use cases. You can check out more about the framework here ([47]).
C.4 CREA’s Multi-Agent Framework Design
@@
C.5 Runtime and Practical Considerations
@@
Appendix D CREA Creative Image Synthesis Examples
D.1 Creative Image Generation
@@ D.1.1 Pre-Generation Planning
D.1.2 Post-Generation Evaluation
After the initial image is generated, it is evaluated in detail by the Art Critic (A4) and the Creative Director (A1) to determine its alignment with the intended creativity blueprint. As shown in Figure 15, the Art Critic applies a multimodal LLM-as-a-Judge framework to assign scores across six creativity principles—Originality, Expressiveness, Aesthetic Appeal, Technical Execution, Unexpected Associations, and Interpretability. Each criterion is scored on a 1–5 scale, and the total is aggregated into a Creativity Index (CI). The Creative Director then reviews these scores in light of the original intent and either approves the image or requests revisions if the CI falls below the creativity threshold. This phase ensures that the image undergoes a rigorous, objective-aligned critique before being finalized.
D.1.3 Self-Enhancement with Optional User-Guidance

### v2

@@ Appendix A Investigating Creativity Types
Appendix B CREA’s Multi-Agentic Architecture
B.1 CREA’s Multi-Agent Framework Design
In the CREA framework, prompts serve as the fundamental coordination interface between specialized agents, guiding the collaborative creative process.Unlike traditional systems that rely on static user prompts, CREA uses modular, role-specific prompting strategies to simulate a human-like creative workflow.These prompts enable agents to reason contextually, communicate asynchronously, and iteratively improve upon the image through multiple refinement cycles.
Our multi-agent framework is composed of five roles: Creative Director, Prompt Architect, Generative Executor, Art Critic, and Refinement Strategist.Each agent is instantiated with a carefully designed system prompt tailored to its unique responsibilities.These prompts are not only essential for aligning agent behavior with creativity principles, but also for facilitating coherent multi-turn conversations between agents. The following subsections describe the prompt structure for each agent, explain its purpose, and highlight how it contributes to the overall creativity loop.For each, we refer to specific examples and prompt templates shown in the supplementary material.
B.2 Creative Director () Prompt
The Creative Director serves as the strategic leader of the agent team.Its prompt equips it not only to interpret the user-provided input (e.g., ‘a couch’) but also to conceptualize a feasible, imaginative, and well-grounded plan for the entire task.More than just an interpreter, the Creative Director performs high-level planning, coordination, and decision-making.It establishes the overall direction of the creative task, specifies whether the objective is image generation or editing, and formulates a comprehensive creativity blueprint.This blueprint includes the visual theme, stylistic and structural constraints, novelty objectives, semantic anchors, and any potential unexpected associations that could add conceptual depth.Once defined, this blueprint is shared with all other agents and acts as the central reference point throughout the iterative process.The Creative Director also participates in post-generation evaluation, arbitrates conflicts in Critic feedback, and decides whether an output is ready for finalization or requires refinement.A high-level system prompt used for the Director, is shown in Template 13.
B.3 Prompt Architect () Prompt
The Prompt Architect operates as the creative translator and synthesis engine of the system.Based on the Creative Director’s blueprint and the creativity template, this agent generates a structured set of prompts to guide the generative process.It first decomposes the blueprint into six contrastive prompts, each aligned with a specific creativity principle—originality, expressiveness, aesthetic appeal, technical execution, unexpected associations, and interpretability.These prompts isolate individual creative dimensions, allowing for focused exploration of each aspect.The template used by to generate Contrastive Prompts is shown in Template 9.Once the contrastive prompts are created, the Prompt Architect uses a Chain-of-Thought fusion process to synthesize them into a single fused high-creativity prompt as shown in Template 10.This fused prompt balances all six principles while preserving semantic clarity and stylistic coherence.Throughout the generation loop, the Prompt Architect also integrates feedback from the Refinement Strategist, modifying or rebalancing prompts as needed to improve underperforming dimensions.This iterative prompting system enables nuanced control over the generation process while ensuring alignment with the blueprint.A high-level system prompt used for the Prompt Architect, is shown in Template 8.
B.4 Generative Executor () Prompt
The Generative Executor is responsible for realizing the creative vision by generating or editing images based on the prompts received from the Prompt Architect.Guided by its system prompt, this agent dynamically interprets the creative instruction and configures the appropriate image generation strategy using T2I models such as Flux or ControlNet.The Executor considers the task type (generation or editing), model-specific parameters like classifier-free guidance or conditioning scale, and any constraints defined by the blueprint.In editing mode, it ensures disentangled transformation by preserving structural elements from the input image while applying stylistic or conceptual changes aligned with the fused prompt.A high-level system prompt used for the Executor, is shown in Template 11.
B.5 Art Critic () Prompt
The Art Critic acts as an autonomous evaluator, responsible for analyzing the generated or edited image and providing structured, multi-dimensional feedback.Its system prompt enables it to assess the output based on the six creativity principles and assign numerical scores from 1 to 5 for each dimension.The Critic also produces rich textual justifications for its ratings and provides an overall creativity score.Beyond evaluation, the Critic serves a pivotal role in steering the refinement process by identifying which dimensions underperform and articulating why.This feedback is critical to both the Creative Director, who decides whether the image meets the creativity threshold, and to the Refinement Strategist, who uses it to guide further improvements.The Critic’s evaluations are grounded in a multi-modal LLM, ensuring assessments are both visually and semantically informed.This agent’s consistent, transparent scoring facilitates traceable progress across iterations. high-level system prompt used for the Critic, is shown in Template 12.
B.6 Refinement Strategist () Prompt
The Refinement Strategist is the system’s adaptive problem solver.Upon receiving the Critic’s evaluation and the current image, this agent analyzes weak areas in the creativity spectrum and formulates a targeted refinement strategy.Its system prompt allows it to propose actionable changes using a formulation that identifies how the existing fused prompt should be adjusted to improve specific dimensions, such as expressiveness or technical execution.The strategist is also tasked with preserving creative coherence; it aligns each suggested change with the original blueprint to avoid drift or over-editing.Its feedback is passed to the Prompt Architect, who updates the prompts accordingly for the next iteration.The Strategist plays a critical role in achieving optimal creativity without compromising conceptual consistency, functioning as the bridge between critique and constructive improvement.A representative example of its prompt behavior and refinement plan is available in Template 14.
B.7 Algorithm
An algorithm summarizing the agentic framework is given in Algorithm 1.
@@ B.7 Algorithm
B.8 Runtime and Practical Considerations
@@
Appendix C CREA Editing and Generation Prompts
We provide the full natural language prompts used to produce the visual results shown in the teaser figure 1 of the main paper.
Examples include both creative image editing (transformations of real-world or AI generated images of objects such as dresses and teapots) and creative image generation (novel interpretations of minimal text inputs such “as "a couch” or “a car”).Each prompt was written to embody multiple creativity principles- e.g., combining unexpected materials, evoking emotional tone, or supporting layered interpretation, resulting in conceptually rich and visually expressive outputs.Figure 11 shows these prompts.
C.1 Creative Image Generation
@@ C.1.1 Pre-Generation Planning
C.1.2 Post-Generation Evaluation
After the initial image is generated, it is evaluated in detail by the Art Critic (A4) and the Creative Director (A1) to determine its alignment with the intended creativity blueprint.As shown in Figure 14, the Art Critic applies a multimodal LLM-as-a-Judge framework to assign scores across six creativity principles: Originality, Expressiveness, Aesthetic Appeal, Technical Execution, Unexpected Associations, and Interpretability.Each criterion is scored on a 1–5 scale, and the total is aggregated into a Creativity Index (CI).The Creative Director then reviews these scores in light of the original intent and either approves the image or requests revisions if the CI falls below the creativity threshold.This phase ensures that the image undergoes a rigorous, objective-aligned critique before being finalized.
C.1.3 Self-Enhancement with Optional User-Guidance